OpenStudy (shamil98):

logx + log(x-2) = log3x The first step is to get everything on one side so.. logx + log(x-2) - log3x = 0 What do I do next?

terenzreignz (terenzreignz):

You have my attention ^_^ Now... all these logs have the same base, yes?

OpenStudy (shamil98):

Yeah, base of x right?

terenzreignz (terenzreignz):

Base of x? I don't think so. In fact, I don't think it would matter what base these logs have, as long as it's the same. Do you see a small number underneath the logs like this: \[\Large \log_{\color{red}b}x\] or not?

OpenStudy (shamil98):

Oh, there's no number.

OpenStudy (shamil98):

\[\large logx + \log(x-2) =\log3x\] that's exactly how it is.

terenzreignz (terenzreignz):

Well then, in that case, this can be solved rather quickly and easily, in two ways, in fact ^_^ First, take advantage of this property: \[\Large \log(MN) = \log(M) + \log(N)\]

OpenStudy (shamil98):

So log x(x-2) = log 3x

terenzreignz (terenzreignz):

I guess you could do that... and then?

terenzreignz (terenzreignz):

Since there is no base (the small number) it is generally assumed that the base is 10. (I think) Regardless, raise 10 to both sides of the equation... like so: \[\Large \log(M) = \log(N)\\\Large 10^{\log(M)}= 10^{\log(N)}\\\Large \ \\\Large M = N\]

OpenStudy (shamil98):

log x(x-2) = log 3x divide by log 3x?.. log x(x-2) / log 3x = 1 x(x-2)/3x = 1 (x-2)/3 = 1 x-2 = 3 x = 5

terenzreignz (terenzreignz):

Answer is correct, but your third line is erroneous ^_^ It should instead be x(x-2)/ 3x = 1 right away, or log [x(x-2)/3x] = 0 But anyway... Here's another way, which, I believe, is less messy :) \[\large \log(x) + \log(x-2)=\log(3x)\]\[\large \log(x)+\log(x-2) = \color{blue}{\log(3) + \log(x)}\] Cancel out log(x) from both sides, we get... \[\Large \log(x-2) = \log(3)\]\[\Large x-2= 3\]\[\Large x = 5 \color{green}{\qquad\checkmark }\]

OpenStudy (shamil98):

Thank you! That makes more sense lol. Forgot about how you can turn log (3x) into log (3) + log (x) .

terenzreignz (terenzreignz):

That doesn't always work, mind you, the way you did it was more...general, save for that tiny error in the middle. Nice work, anyway. ^_^